OpenStudy (anonymous):

write 3^-3 with no negative exponents

OpenStudy (anonymous):

\(b^{-n}=\frac{1}{b^n}\)

OpenStudy (kainui):

Since we know stuff like: \[\frac{ 5^7 }{ 5^3 } =5^4\] Then you can see how 1/5^3 is the same as 5^-3